OpenStudy (anonymous):

what do m= for the equation 5(1+4m)=2(3+10m)

OpenStudy (anonymous):

can you solve this problem

OpenStudy (anonymous):

is there a solution

OpenStudy (anonymous):

Distribute my good sir. 5(1+4m)=2(3+10m) 5+20m=6+20m. 20m=1+20m 0=1. This obviously doesn't work. The equation has no solution and is false.